A water sprinkler has a radius of 6 feet. How much area does the water sprinkler cover? Use 3.14 for and round your answer to the nearest tenth.


Answer by rfer(16322)

a=pir^2
a=3.14*6^2
a=3.14*36
a=113.0 sq ft